A delivery car had a first cost of $36,000, an annual operating cost of $17,000, and an estimated $3500 salvage value after its 6-year life. Due to an economic slowdown, the car will be retained for only 3 years and must be sold now as a used vehicle. At an interest rate of 9% per year, what must the market value of the used vehicle be in order for its AW value to be the same as the AW if it had been kept for its full life cycle?

The market value of the used vehicle is determined to be ___________$  .
Answer #1

AW of machine 6 yrs = -36000 * (A/P,9%,6) - 17000 + 3500 * (A/F,9%,6)

= -36000 *0.222920 - 17000 + 3500 *0.132920

= -24559.90

Let Salvage value be S after 3 yrs

AW of machine 3 yrs = -36000 * (A/P,9%,3) - 17000 + S * (A/F,9%,3)

= -36000 *0.395055 - 17000 + S * 0.305055

= -31221.98 + S * 0.305055

As per given condition

-31221.98 + S * 0.305055 = -24559.90

S * 0.305055 = -24559.90 + 31221.98 = 6662.08

S = 6662.08 / 0.305055 = 21838.96 ~ 21839 (Nearest Dollar)